# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ Abstract
In this paper, we explore a class of constrained saddle point problems with a bilevel structure, where the upper-level objective function is nonconvex-concave and smooth subject to a strongly convex lower-level objective function. This class of problems finds wide applicability in machine learning, including robust multi-task learning, adversarial learning, and robust meta-learning. While some studies have focused on simpler formulations, e.g., when the upper-level objective function is linear in the maximization component, there remains a significant gap in developing efficient projection-free and projection-based algorithms with theoretical guarantees for more general settings. To bridge this gap, we propose efficient single-loop one-sided projection-free, and fully projection-based primal-dual methods. By leveraging regularization and nested approximation techniques, we initially devise a bilevel primal-dual one-sided projection-free algorithm, requiring $\mathcal{O}(\epsilon^{-4})$ iterations to attain an $\epsilon$-stationary point. Subsequently, we develop a bilevel primal-dual fully projected algorithm, capable of achieving an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations. To the best of our knowledge, our proposed algorithms represent among the first methods for solving general non-bilinear saddle point problems with a bilevel structure.
@@ 1 Introduction
Bilevel optimization is an important class of optimization problems where one optimization problem is nested within another. Many learning and inference problems take this form with two levels of hierarchy. On the other hand, Saddle Point (SP), also known as min-max, optimization problems have received a lot of attention recently due to their versatile formulation covering a broad class of optimization problems, e.g., convex optimization with nonlinear conic constraints, which itself includes LP, quadratic programming (QP), quadratically constrained quadratic programming (QCQP), second-order cone programming (SOCP), and semidefinite programming (SDP) as its subclasses. From a broader perspective, in this paper, we study a saddle point problem with bilevel structure (SPB) of the following form
@@
where $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\to\mathbb{R}$ is a continuously differentiable function such that $\Phi(x,\theta,\cdot)$ is concave for any given $x\in\mathcal{X}$ and $\theta\in\mathbb{R}^{m}$; $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{d}$ are convex and compact sets; moreover, $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a continuously differentiable and strongly convex in $\theta$. This problem simplifies to bilevel optimization when the upper-level objective function $\Phi$ is separable in the maximization variable and it encompasses saddle point problems when there is no lower-level problem, i.e., $\Phi$ is independent of the second argument. The problem formulated in (1) has a wide range of applications including robust multi-task learning [39], robust meta-learning [19, 24], adversarial learning [57], representation learning and hyperparameter optimization [20], multi-task deep AUC maximization, and multi-task deep partial AUC maximization [25].
@@ 1.1 Literature Review
Bilevel Optimization. Researchers have proposed numerous algorithms for solving bilevel optimization problems since its initial introduction by Bracken and McGill in 1973 [5]. These methods include but are not limited to, constraint problem [40, 49] and unconstrained [26, 44] approaches. In recent years, there has been a growing research focus on applying gradient-based settings to solve bilevel problems, incorporating techniques such as implicit differentiation [12, 15, 17] and iterative differentiation [13, 18, 38, 47]. In the literature, several single and double-loop approaches have been developed to address unconstrained bilevel optimization problems [10, 28, 30, 32, 35, 54]. There are fewer studies focusing on constrained bilevel optimization problems, particularly those that consider a class of bilevel optimization problems with an upper-level objective function over a constraint set. Researchers have proposed a number of methods, including projection-based [15, 23], and projection-free [1, 2] algorithms for solving constrained bilevel optimization problems. Among the projection-based methods, specifically, the authors in [23] proposes a two-timescale stochastic approximation (TTSA) algorithm for solving a bilevel optimization problem. In this scenario, the upper-level problem is constrained and has a nonconvex objective function, while the lower-level problem is unconstrained and strongly convex. By employing a projection onto the set $\mathcal{X}$ at each iteration, their algorithm achieves a complexity of $\mathcal{O}(\epsilon^{-5/2})$ in this setting. In contrast, there is an increasing interest in using projection-free methods, which aim to reduce the computational cost associated with projections at each iteration through the utilization of a linear solver for solving the constrained setting. Recently, researchers in [1] study a single-loop projection-free method called inexact bilevel conditional gradient (IBCG) for solving a bilevel optimization problem with a smooth and constrained upper-level objective function and a strongly convex lower-level problem. Their method converges to an $\epsilon$-optimal solution with a complexity of $\mathcal{O}(\epsilon^{-2})$ when the upper-level objective function is non-convex. The convergence rates for both the TTSA and IBCG algorithms match those of their single-level counterparts.
@@
Saddle Point (SP) Problem. In the context of saddle point (SP) problems, some research has focused on the convex-concave setting [6, 21, 41, 43], while there is an increasing number of studies concentrating on developing algorithms for nonconvex-concave setting [4, 36]. In the nonconvex-concave setting, primarily two types of algorithms, namely projected gradient-based algorithms and projection-free-based algorithms, are studied. The projected gradient-based algorithms can be applied in various methods, including but not limited to adding a momentum term [21, 52], accelerated gradient methods [22, 29], gradient descent-ascent [11, 55], multi-loop algorithms [37, 42]. To ensure a guaranteed convergence rate, numerous methods, including smoothing techniques [56] and acceleration methods [42], have been investigated. Many of these methods are based on solving regularized subproblems inexactly, leading to a multi-loop approach that can present implementation challenges. To overcome such complexities, there has been a growing preference for single-loop primal-dual methods. However, fewer research studies cover single-loop algorithms for the nonconvex-concave setting. In [53], the authors introduce a unified single-loop alternating gradient projection (AGP) algorithm designed for solving smooth nonconvex-(strongly) concave problems. Their algorithm can find an $\epsilon$-stationary point of the objective function, achieving a complexity of $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ under the nonconvex-strongly concave and nonconvex-concave settings, respectively.In the optimization literature, limitations in projection-based techniques have prompted researchers to explore novel and efficient projection-free approaches [31, 33, 34, 46, 51]. In the context of saddle point (SP) problems, a limited amount of research has been conducted on projection-free SP problems [4, 9, 16]. In particular, [4] investigates single-loop fully and one-sided projection-free algorithms designed to solve a saddle point problem with a smooth nonconvex-concave objective function. Their proposed method can identify an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-6})$ and $\mathcal{O}(\epsilon^{-4})$ for fully and one-sided projection-free methods, respectively.
@@
Saddle Point Problem with Bilevel Structure (SPB). Within the existing literature, the focus on the SPB problem, as presented in Problem (1), is limited. [25] proposes two simple, single-loop, single-timescale stochastic methods with randomized block-sampling to solve the general form of Problem (1). This is done under the assumption that the upper-level objective function $\Phi(x,\theta^{*}(x),y)$ is a smooth, unconstrained, nonconvex strongly-concave function, and the lower-level objective function $g(x,\theta)$ is smooth and strongly-convex. Theoretically, they show that the algorithms converge to an $\epsilon$-stationary point with complexity $\mathcal{O}(\epsilon^{-4})$ under a general unbiased stochastic oracle model. In these approaches, Stochastic Gradient Descent (SGD) is utilized to update selected $\theta_{i}$ for their respective lower-level problems, Stochastic Gradient Ascent (SGA) is employed for updating the selected $y_{i}$, and a momentum update is applied to the primal variable $x$ based on the sampled $y_{i}$ and $\theta_{i}$. Contrastingly, the upper-level objective function in our problem is nonconvex-concave over a constraint set. Additionally, we propose a projection-free method to alleviate the computational cost associated with the application of projection-based methods. In another study, [19] employs a minimax bilevel optimization problem to formulate the task-robust meta-learning problem. They consider a minimax bilevel optimization problem of the form $\min_{x\in\mathcal{X}}\max_{i\in[n]}\Phi_{i}(x,\theta^{*}(x))$ subject to $\theta^{*}(x)\in\operatorname*{\arg\!\min}_{y_{i}\in\mathcal{Y}_{i}=\mathbb{R}^{d_{y_{i}}}}g_{i}(x,y_{i}),\forall i\in[n]$, a special case of Problem (1), assuming that the objective function $\Phi$ is linear in $y$. Their proposed algorithm, TaRo-BOBA (Task Robust Bilevel Optimization Bi-timescale Algorithm), is a single-loop, two-timescale, stochastic algorithm based on gradient descent-ascent to solve their minimax bilevel optimization problem. They assume that their upper-level problem is constraint nonconvex-concave, and the lower-level objective function is strongly-convex and unconstrained. According to theoretical studies, TaRo-BOBA converges to the first-order stationary point at a rate of $\mathcal{O}(\sqrt{n}\epsilon^{-5})$, where $n$ denotes the number of tasks. In a recent study [20], a minimax multi-objective bilevel optimization problem is explored, with applications in robust machine learning. They propose a single-loop Multi-Objective Robust Bilevel Two-timescale optimization algorithm, MORBiT, in a similar setting and convergence rate as discussed in [19]. In [24], the authors propose a single-loop, single-timescale stochastic algorithm using a stochastic estimator based on moving average (SEMA) to solve a nonconvex minimax bilevel optimization problem. Specifically, they investigate a problem where the upper-level objective function is nonconvex and strongly-concave, while the lower-level objective is considered strongly-convex. They theoretically demonstrate that their method can achieve an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-4})$ iterations, which yields better results compared to a similar study in [19].
There are a limited number of studies considering the SPB problem. Existing methods primarily focus on projection-based algorithms, and, to the best of our knowledge, there is no exploration of any projection-based and projection-free algorithm alternatives in this setup. Therefore, in this paper, we concentrate on developing both projection-free and projection-based methods for solving Problem (1).
@@ 1.2 Contribution
In this paper, we consider a class of constrained saddle point problems with a bilevel structure, where the upper-level objective function is nonconvex-concave and smooth over a compact and convex constraint set, subject to a strongly convex lower-level objective function. This study extends the current literature, which has primarily focused on using projection-based methods to solve nonconvex strongly-concave upper-level objective functions [24, 25], or where their upper-level objective function is linear in the dual variable [19, 20]. Our motivation stems from the need to develop efficient algorithms both projection-free and projection-based with theoretical guarantees, addressing the gap in methods for solving general non-bilinear SPB problem.
Utilizing regularization and nested approximation, we first develop a single-loop, one-sided projection-free algorithm. In the upper-level objective function, this algorithm uses a linear minimization oracle (LMO) and a regularized projection gradient ascent to find directions for primal and dual steps, respectively. Next, we develop a regularized single-loop projected gradient-based algorithm that uses a projection oracle (PO) to update the primal and dual steps of the upper-level objective function. In both algorithms, we incorporate a one-step gradient-type update to estimate the lower-level optimal solution and its derivative.
Our main theoretical guarantees for the proposed methods are summarized as follows:
  • The proposed bilevel primal-dual one-sided projection-free method achieves an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-4})$ iterations.
  • The proposed bilevel primal-dual fully projected method requires $\mathcal{O}(\epsilon^{-5})$ iterations to find an $\epsilon$-stationary point.
@@ 1.3 Motivating Examples
There are some applications of SPB formulation in Problem (1), including machine learning problems such as adversarial learning [57], representation learning and hyperparameter optimization [20], robust multi-task learning [39], robust meta-learning [19, 24], multi-task deep AUC maximization, and multi-task deep partial AUC maximization [25]. Next, we describe two examples in detail.
Adversarial Learning: With the advancement of technology, it is crucial to enhance the robustness and security of machine learning models against adversarial attacks. Adversarial learning aids in defense against attacks and enhances generalization by training models to withstand adversarial examples. Recently, in [57], the authors propose a bilevel form of adversarial training (AT), representing a unified bilevel optimization problem with a discriminator model in the upper-level and a generator model in the lower-level. Specifically, their proposed algorithm is interpreted as a bilevel optimization with a linearized lower-level problem. However, their formulation assumes knowledge of the data distribution, which is problematic in real-world applications.
To address this, one approach is to utilize distributionally robust optimization (DRO) to develop models resilient to variations in data distribution. Instead of relying on a specific assumed distribution, DRO considers a set of possible distributions, represented by an uncertainty set. This approach ensures that the solution performs well across a range of possible distributions, enhancing robustness to uncertainties in the data. Therefore, by combining the ideas of AT in bilevel optimization and DRO, we can present the problem generally as follows:
$\min_{x\in\mathcal{X}}\max_{P\in\mathcal{P}}\mathbb{E}_{P}[\hat{\ell}(x,\theta^{*}(x);\mathcal{D}^{tr})]$
@@
where $\hat{\ell}$ and $\ell_{atk}$ represent the training loss function and the attack objective function, respectively. $\mathcal{X}$ denotes a convex set, and $\mathcal{P}$ signifies an uncertainty set. Furthermore, we can revisit Problem (1.3) by utilizing sample average approximation to approximate the expectation, and by linearizing $\ell_{atk}$, we can formulate it similarly to Problem (1) as:
$\min_{x\in\mathcal{X}}\max_{P\in\mathcal{P}}\frac{1}{n}\sum_{i=1}^{n}P_{i}\hat{\ell}(x,\theta^{*}(x);\mathcal{D}^{tr}_{i})$
@@
where $P_{i}$ is the probability distribution for data point $\mathcal{D}^{tr}_{i}$ for $i\in[N]\triangleq\{1,2,…,n\}$, $\overline{\theta}$ is given, and $\mathcal{R}(\theta)$ is a regularization term such as $L_{2}$ regularization.
Robust Multi-task Learning: Consider a multi-task learning (MTL) problem where the goal is to train a model on multiple related tasks simultaneously, aiming to create specialized models for each task that perform better compared to training separate models for each task. The common approach to train the model for MTL is to minimize the average of all task losses. However, a drawback of this method is that all tasks are considered equally, while it is possible that the distribution of tasks is heterogeneous, causing a skewness in the performance of the acquired model toward certain tasks.
To overcome this problem, one approach is to minimize the worst task loss of the model, making the attained model robust against the distribution differences among tasks. This can be viewed as an SPB problem, wherein the upper level (UL), we have a saddle point problem, and in the lower level (LL), we have a minimization problem parameterized by the solution of the UL problem. This formulation has arisen in many examples. In task robust meta-learning problem [19, 24], a similar structure is utilized, where the primal variable is considered as a meta-learner, and the shared parameter is used as the task-specific learners in the adaptation phase. In the robust representation learning problem [20], a similar formulation is employed, where a shared representation network is regarded as the upper-level (UL) problem, and the lower-level (LL) problem is viewed as a per-task linear model parameterized with the solution of the UL problem, aiming to learn a shared representation and per-task models so that each task generalizes effectively and achieves robust performance across tasks. Generally, this problem can be expressed as follows:
$\min_{x\in\mathcal{X}}\max_{y\in\Delta_{N}}\sum_{i=1}^{N}y_{i}\ell_{i}(x,\theta_{i}^{*}(x);\mathcal{D}_{i}^{val})$
@@
where $\mathcal{X}$ is a convex set (e.g. $\|x\|_{1}\leq r$), $[N]\triangleq\{1,2,…,n\}$ is the index of each learning task, and $\Delta_{N}$ is the simplex $\Delta_{N}\subset\mathbb{R}^{n}=\{y\in\mathbb{R}^{n}|y_{1}+…+y_{n}=1,y_{i}\geq 0\}$. In this setup, the lower-level and upper-level objectives correspond to a learning task $i\in[n]$ with its training and validation dataset pair $\mathcal{D}_{i}^{tr},\mathcal{D}_{i}^{val}$, respectively. In Problem (1.3), $\mathcal{R}(\theta)$ is a regularization term such as $L_{2}$ regularization, utilized to ensure the strong convexity of the lower-level problem.
An alternative formulation for robust MTL can be considered as a weight selection for MTL-DRO. Several research [8, 27, 45, 50] consider the concept of sample reweighting, which can be formulated as a bilevel optimization problem. In another study, utilizing distributionally robust optimization (DRO), [39] proposes a robust multi-task learning (MTL) problem in a single-level saddle point structure where the worst-task loss of the model is minimized at each step. Considering both ideas together, the problem can be formed as an SPB as follows:
$\min_{x\in\Delta_{N}}\max_{y\in\mathcal{U}}\sum_{i=1}^{N}y_{i}\ell_{i}(\theta^{*}(x);\mathcal{D}_{i}^{val})$
[…]
where $\Delta_{N}$ is the simplex $\Delta_{N}\subset\mathbb{R}^{n}=\{x\in\mathbb{R}^{n}|x_{1}+…+x_{n}=1,x_{i}\geq 0\}$, $\mathcal{U}$ is an uncertainty set, and $\mathcal{R}(\theta)$ denotes a regularization term for ensuring strong convexity of the lower-level problem.

### v3

@@ Abstract
In this paper, we explore a broad class of constrained saddle point problems with a bilevel structure, wherein the upper-level objective function is nonconvex-concave and smooth over compact and convex constraint sets, subject to a strongly convex lower-level objective function. This class of problems finds wide applicability in machine learning, encompassing robust multi-task learning, adversarial learning, and robust meta-learning. Our study extends the current literature in two main directions: (i) We consider a more general setting where the upper-level function is not necessarily strongly concave or linear in the maximization variable. (ii) While existing methods for solving saddle point problems with a bilevel structure are projected-based algorithms, we propose a one-sided projection-free method employing a linear minimization oracle. Specifically, by utilizing regularization and nested approximation techniques, we introduce a novel single-loop one-sided projection-free algorithm, requiring $\mathcal{O}(\epsilon^{-4})$ iterations to attain an $\epsilon$-stationary solution, moreover, when the objective function in the upper-level is linear in the maximization component, our result improve to $\mathcal{O}(\epsilon^{-3})$. Subsequently, we develop an efficient single-loop fully projected gradient-based algorithm capable of achieving an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-5})$ iterations. This result improves to $\mathcal{O}(\epsilon^{-4})$ when the upper-level objective function is strongly concave in the maximization component. Finally, we tested our proposed methods against the state-of-the-art algorithms for solving a robust multi-task regression problem to showcase the superiority of our algorithms.
@@ 1 Introduction
Bilevel optimization is an important class of optimization problems where one optimization problem is nested within another. Many learning and inference problems take this form with two levels of hierarchy. On the other hand, saddle point (SP) problems, also known as min-max optimization, have received a lot of attention recently due to their versatile formulation covering a broad class of optimization problems, e.g., convex optimization with nonlinear conic constraints, which itself includes linear programming (LP), quadratic programming (QP), quadratically constrained quadratic programming (QCQP), second-order cone programming (SOCP), and semidefinite programming (SDP) as its subclasses. From a broader perspective, in this paper, we study a saddle point problem with a bilevel structure (SPB) of the following form
@@
where $\Phi:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{d}\to\mathbb{R}$ is a continuously differentiable function such that $\Phi(x,\theta,\cdot)$ is concave for any given $x\in\mathcal{X}$ and $\theta\in\mathbb{R}^{m}$; $\mathcal{X}\subset\mathbb{R}^{n}$ and $\mathcal{Y}\subset\mathbb{R}^{d}$ are convex and compact sets; moreover, $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a continuously differentiable and strongly convex in $\theta$. This problem simplifies to bilevel optimization when the upper-level objective function $\Phi$ is separable in the maximization variable and it encompasses saddle point problems when there is no lower-level problem, i.e., $\Phi$ is independent of the second argument. The problem in (1) has a wide range of applications including robust multi-task learning [36], robust meta-learning [20], adversarial learning [48], representation learning and hyperparameter optimization [21], multi-task deep AUC maximization, and multi-task deep partial AUC maximization [24].
@@ 1.1 Literature Review
Bilevel Optimization.
Researchers have proposed numerous algorithms for solving bilevel optimization problems since its initial introduction by Bracken and McGill in 1973 [5]. In recent years, there has been growing research focus on applying gradient-based methods for solving bilevel problems, incorporating techniques such as implicit differentiation [12, 15, 18] and iterative differentiation [13, 19, 34, 41]. The majority of existing studies proving convergence rate guarantees have focused on unconstrained scenario [11, 26, 27, 29, 32, 46], while fewer studies are considering constrained bilevel optimization problems, particularly those with constraints in the upper-level. In particular, researchers have proposed several methods, including projection-based [15, 23] and projection-free [1, 2] algorithms for various settings to achieve complexity guarantees matching the single-level counterpart.
Saddle Point (SP) Problem.
@@
In the context of saddle point (SP) problems, there have been many studies on the convex-concave setting [7, 22, 38, 40], while there is an increasing number of studies concentrating on developing algorithms for nonconvex-concave setting [4, 33, 39, 47]. In the nonconvex-concave setting, various approaches including smoothing [47] and acceleration [39] techniques have been porposed. Most of these methods are based on solving regularized subproblems inexactly, leading to a multi-loop algorithm that can present implementation challenges. To overcome such complexities, there has been a growing interest in single-loop primal-dual methods. In particular, the authors in [45] introduce a unified single-loop alternating gradient projection (AGP) algorithm designed for solving smooth nonconvex-(strongly) concave problems. Their algorithm can find an $\epsilon$-stationary solution with a complexity of $\mathcal{O}(\epsilon^{-2})$ and $\mathcal{O}(\epsilon^{-4})$ under the nonconvex-strongly concave and nonconvex-concave settings, respectively. In the optimization literature, limitations in projection-based techniques have prompted researchers to explore efficient projection-free approaches [14, 28, 30, 31]. In the context of saddle point (SP) problems, a limited number of studies have been conducted on projection-free SP problems [4, 10, 16]. In particular, authors in [4] propose fully and one-sided projection-free algorithms designed to solve an SP problem with a smooth nonconvex-concave objective function. Their proposed method can achieve an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-6})$ and $\mathcal{O}(\epsilon^{-4})$ iterations for fully and one-sided projection-free methods, respectively.
Saddle Point Problem with Bilevel Structure.
@@
Within the existing literature, the focus on the SPB problem, as presented in problem (1), is very limited. In [24] two simple, single-loop, single-timescale stochastic methods with randomized block-sampling have been proposed to solve (1). Considering an unconstrained setting and assuming that $\Phi$ is smooth, $\Phi(x,\theta,\cdot)$ is strongly concave for any $x,\theta$, and $g(x,\cdot)$ is strongly convex for any $x$, they show that their proposed algorithm converges to an $\epsilon$-stationary point with a complexity of $\mathcal{O}(\epsilon^{-4})$ under a general unbiased stochastic oracle model. In another study, [20] motivated by a robust meta-learning problem, consider a bilevel optimization with minimax upper-level objective function, i.e., $\min_{x\in\mathcal{X}}\max_{i=1}^{n}f_{i}(x,\theta_{i}^{*}(x))$. This problem can be reformulated as a special case of (1) by letting $\Phi(x,\theta,y)=\sum_{i=1}^{n}y_{i}f_{i}(x,\theta_{i})$, i.e., $\Phi$ is assumed to be linear in $y$. Assuming that $\Phi$ is smooth, their proposed algorithm can achieve a complexity of $\mathcal{O}(\epsilon^{-5})$. In recent studies [20, 21], a minimax multi-objective bilevel optimization problem with applications in robust machine learning is explored. They propose a single-loop Multi-Objective Robust Bilevel Two-timescale optimization (MORBiT) algorithm for solving (1) in a deterministic regime when $\Phi$ is linear in $y$. Their algorithm converges to an $\epsilon$-stationary solution at a rate of $\mathcal{O}(\epsilon^{-5})$.
@@ 1.2 Contribution
In this paper, we consider a broad class of constrained SP problems with a bilevel structure, where the upper-level objective function is smooth over compact and convex constraint sets, subject to a strongly convex lower-level objective function. While the existing methods for solving (1) focus on special cases of this problem when either $\Phi$ is linear or strongly concave in $y$, we consider a more general setting where the $\Phi(x,\theta,\cdot)$ is a concave function for any $x,\theta$. Utilizing regularization and nested approximation, we propose novel single-loop inexact bilevel primal-dual algorithms featuring only two matrix-vector products at each iteration. To contend with the challenge of projection onto the constraint set $\mathcal{X}$, we propose a one-sided projection-free method utilizing a linear minimization oracle. Our main theoretical guarantees for the proposed methods are summarized as follows:
  • The proposed Inexact Bilevel Regularized Primal-dual One-sided Projection-free (i-BRPD:OPF) method achieves an $\epsilon$-stationary solution within $\mathcal{O}(\epsilon^{-4})$ iterations. This complexity result improves to $\mathcal{O}(\epsilon^{-3})$ when $\Phi$ is linear in $y$.
  • The proposed Inexact Bilevel Regularized Primal-dual Fully Projected (i-BRPD:FP) method requires $\mathcal{O}(\epsilon^{-5})$ iterations to find an $\epsilon$-stationary point. This complexity result improves to $\mathcal{O}(\epsilon^{-4})$ when $\Phi$ is strongly concave in $y$.
To the best of our knowledge, the proposed methods and their convergence guarantee are the first to appear in the literature for the considered setting.
@@ 1.3 Motivating Examples
The SPB problem in (1) presents a versatile formulation covering various applications arising in machine learning such as adversarial learning [48], representation learning and hyperparameter optimization [21], robust multi-task learning [36], robust meta-learning [20], and multi-task deep AUC maximization [24]. Next, we describe two examples in detail.
Robust Multi-task Learning: Multi-task learning (MTL) is a machine learning approach that focuses on jointly learning multiple related tasks [6]. The main objective is to share knowledge across these tasks, allowing them to benefit from each other’s insights. More specifically, consider a set of $T$ tasks denoted by $\{\mathcal{T}_{i}\}_{i=1}^{T}$, each equipped with a training dataset $\{\mathcal{D}_{i}^{tr}\}_{i=1}^{T}$ containing features residing in the same space. In the context of these tasks, we employ task-specific loss functions $\{\ell_{i}(x,\theta_{i},\mathcal{D}_{i}^{tr})\}_{i=1}^{T}$ where $\theta_{i}\in\mathbb{R}^{m}$ represents task-specific model parameters, and $x$ is the shared parameter. Although many formulations and loss functions have been proposed in the MTL literature, they generally yield a weighted empirical risk minimization $\min_{\theta=[\theta_{i}]_{i=1}^{T}}\sum_{i=1}^{T}y_{i}\ell_{i}(x,\theta_{i};\mathcal{D}_{i}^{tr})$, where the weights $\{y_{i}\}_{i=1}^{T}$ play a crucial role in the performance of the model as more favor is given to the tasks with higher weights. Therefore, it is natural to consider learning the weights through a robust optimization formulation [9, 43] and learning the shared parameter $x$ through the lens of bilevel optimization [20]. This problem can be uniformly formulated as
$\min_{x\in\mathcal{X}}\max_{y\in\mathcal{U}}\sum_{i=1}^{T}y_{i}\ell_{i}(x,\theta_{i}^{*}(x);\mathcal{D}_{i}^{val})$
@@
where $\mathcal{U}\subset\mathbb{R}^{T}$ is the weight constraint set, e.g., $\mathcal{U}=\Delta_{T}$ and $\Delta_{T}$ denotes the simplex set defined as $\Delta_{T}\triangleq\{y\in\mathbb{R}^{T}_{+}|\sum_{i=1}^{T}y_{i}=1\}$, $\mathcal{D}_{i}^{val}$ is the validation dataset for task $i$, and $\mathcal{R}:\mathbb{R}^{m}\to\mathbb{R}$ is a regularization function, e.g., $\mathcal{R}(\theta_{i})=\frac{\rho}{2}\left\|\theta_{i}\right\|^{2}$ for some $\rho>0$.
Adversarial Learning: As safeguarding machine learning models against adversarial attacks becomes increasingly vital, adversarial learning emerges as a pivotal strategy in fortifying these models. By exposing models to adversarial examples during training, adversarial training (AT) [17, 35] significantly enhances their robustness against potential attacks. Recently, in [48], the authors propose a bilevel formulation of the AT, wherein they aim to minimize the expectation of the training loss function at the upper-level formulated as $\min_{x\in\mathcal{X}}\mathbb{E}_{\mathcal{D}^{tr}\sim\mathcal{P}}[\ell(x,\theta^{*}(x);\mathcal{D}^{tr})]$ while minimizing the attack loss function at the lower-level formulated as $\theta^{*}(x)\in\operatorname*{\arg\!\min}_{\theta}\ell_{atk}(x,\theta)$. In this formulation, the upper-level involves an expectation that assumes knowledge of the data distribution. However, in real-world applications, such knowledge may not be available. To address this, one standard approach is to utilize the distributionally robust optimization (DRO) [37] technique. Instead of relying on a specific assumed distribution, DRO considers a set of possible distributions, represented by an uncertainty set. Therefore, integrating a DRO formulation within bilevel AT we can formulate the following robust adversarial learning problem:
$\min_{x\in\mathcal{X}}\max_{p=[p_{i}]_{i=1}^{N}\in\mathcal{U}}~{}\sum_{i=1}^{N}p_{i}{\ell}(x,\theta^{*}(x);\mathcal{D}^{tr}_{i})$
@@
where $\mathcal{U}\subset\mathbb{R}^{N}$ denotes the uncertainty set. Furthermore, when the uncertainty set is represented by a divergence measure, e.g., $\mathcal{U}=\{p\in\Delta_{N}\mid V(p,\frac{1}{N}\mathbf{1}_{N})\leq\rho\}$ for some divergence metric $V$ and $\rho>0$, then projection onto $\mathcal{U}$ is computationally expensive. Therefore, using the Lagrangian method and introducing variable $\lambda\in\mathbb{R}_{+}$ we can relax the nonlinear constraint in $\mathcal{U}$ to equivalently represent the problem as
$\min_{\begin{subarray}{c}x\in\mathcal{X}\\ \lambda\geq 0\end{subarray}}\max_{p\in\Delta_{N}}~{}\sum_{i=1}^{N}p_{i}{\ell}(x,\theta^{*}(x);\mathcal{D}^{tr}_{i})]-\lambda\big{(}V(p,\tfrac{1}{N}\mathbf{1}_{N})-\rho\big{)}$ (2)
@@
Indeed, one can easily verify that (2) is a special case of (1). Note that the upper-level objective function in (2) is neither linear in the maximization component $p$ nor strongly concave.